Prove that for all n ≥ 1, nn ≥ n!.
Answer #1
for n=1:
---------
LHS = 1^1 = 1
RHS = 1! = 1
so, LHS >= RHS for n=1

let's assume LHS >= RHS for n=k:
---------------------------------
k^k >= k!

for n=k+1:
-----------
LHS = (k+1)^(k+1)
= (k+1)^k * (k+1)
>= k^k * (k+1)
>= k! * (k+1)
>= (k+1)!
>= RHS

so, LHS >= RHS for n=k+1, given that LHS >= RHS for n=k.
hence proved by induction.
Answer #2

for n=1:

LHS = 1^1 = 1,RHS = 1! = 1 

so, LHS >= RHS for n=1 

let's assume LHS >= RHS for n=k: 

k^k >= k!

 for n=k+1:

 LHS = (k+1)^(k+1) = (k+1)^k * (k+1) >= k^k * (k+1) >= k! * (k+1) >= (k+1)! >= RHS 

so, LHS >= RHS for n=k+1


answered by: Shivani Sharma